what mass of Cobalt contains the same number of atoms as 57.0 g of flourine
1
Expert's answer
2016-10-25T14:58:10-0400
Number of moles is the measure of atom number (It is defined as the amount of a chemical substance that contains as many elementary entities, e.g., atoms, molecules, ions, electrons, or photons, as there are atoms in 12 grams of carbon-12 (12C), the isotope of carbon with relative atomic mass 12 by definition).

n = m/M
M (Co) = 58.9 g/mol
M (F) = 18.9 g/mol
n (F) = 57.0 / 18.9 = 3.0 moles
m (Co) = 3.0 · 58.9 = 177.6 g